OpenStudy (anonymous):

three more than 8 times a number is -29.find the number

OpenStudy (anonymous):

8x+3 = -29

OpenStudy (anonymous):

8n+3=-29 8n=-29-3 8n=-32 n= -32/8 n=-4

OpenStudy (anonymous):

8x = -32 x=-4

OpenStudy (anonymous):

ThankYouu (: